The sum of the digits of a two digit number is four more than the excess of the units digit over the tens digit.if the product of the digit is 16 .find the number
Answer by lwsshak3(11628)

The sum of the digits of a two digit number is four more than the excess of the units digit over the tens digit.if the product of the digit is 16 .find the number
**
let u=units digit
let t=tens digit
u*t=16
u=16/t
..
u+t=(u-t)+4
2t=4
t=2
u=16/t=8
number: 28